OpenStudy (anonymous):

Given the linear equation y=1/3x - 3, find the y coordinates of the point (-6), (-3), and (6).

OpenStudy (anonymous):

substitute x with the points

OpenStudy (anonymous):

ok it would be -6/3 = 2 ?

OpenStudy (anonymous):

y=1/3(-6)-3 y=-2-3 y=-5 so the first point would be at (-6,-5)

OpenStudy (anonymous):

do the same for the other points y=1/3(-3)-3 and y=1/3(6)-3

OpenStudy (anonymous):

so my second points would be (-3,0)

OpenStudy (anonymous):

well if you do y=1/3(-3)-3 it becomes y=(-3)/3-3 y=-1-3=-4 so the second point would be (-3,-4)

OpenStudy (anonymous):

ok i see

OpenStudy (anonymous):

so this one would be y = 1/3 (6) -3 y = 2 -3 y = -1 (-1, -3)

OpenStudy (anonymous):

nono you're solving for the y-coordinate given the x-coordinate, so the answer would be (6,-1)

OpenStudy (anonymous):

so the 6 they give you is the x coordinate and you just solved for the y coordinate which is -1 and then you just put them together

OpenStudy (anonymous):

oh ok i got it now. thank you so much

OpenStudy (anonymous):

np